The rate of a certain reaction was studied at various temperatures. The table shows temperature (T) and rate constant (k) data collected during the experiments. Plot the data, and then answer the following questions.

What is value of the activation energy, Ea, for this reaction?

What is the value of the pre-exponential factor (sometimes called the frequency factor), A, for this reaction?

The rate of a certain reaction was studied at vari
Answer #1

For this we have to first plot 1/T versus ln k then we have slope = -Ea/R

See plot

slope = -13299

Ea/R = 13299

Ea = 13299 x8.314 = 110.6 kJ/mol

The intercept gives the value of = ln(A)

26.046 = ln(A)

A = e26.046

A = 2.04 x 1011s-1z -2 -4 -6 у:-13299( + 26.046 -8 0.0015 0.0017 0.0019 0.0021 0.0023 0.0025 0.0027 1/Т (К-1)
Answer #2

image.png

source: Macmillan
answered by: RawReciever